OpenStudy (anonymous):

Find F inverse prime at a=4 F(x)=2x^2+3x^2+7x+4 I have the Derivative of the inverse already I just want to know how you know f'(4)=0..

OpenStudy (anonymous):

The Derivative of the inverse is y'=1/(6y^2+6y+7)

OpenStudy (anonymous):

First question, in the inverse replace all x's with 4 and solve... what did you get for the inverse?

OpenStudy (anonymous):

2x^3 + 3x^2 + 7x + 4 = 4 2x^3 + 3x^2 + 7x = 0 x(2x^2 + 3x + 7) = 0 The quadratic in the brackets has no solutions (discriminant is negative), so only solution is x = 0.

OpenStudy (anonymous):

I did implicate differentiation to get the derivative of the inverse function so I just replaced all the x with y... But If I replace all the x's with 4 it becomes a huge number not 0 specifically 208.

OpenStudy (anonymous):

Oh that makes sense okay so would I do it like that on all problems? Just make the problem = to whatever a =?

OpenStudy (anonymous):

So (f^-1)(4) = 0 f ' (x) = 6x^2 + 6x + 7 f ' (0) = 7 1 / f ' (0) = 1/7

OpenStudy (anonymous):

Thank you! What happens if x = 2 numbers what do I do then?

OpenStudy (anonymous):

find the derivative and replace x with 2 then take the inverse

OpenStudy (anonymous):

Thanks! Would you be able to help me with one more problem?

OpenStudy (anonymous):

So let a = f(x). Then: (f^-1) ' (a) = 1 / f ' [(f^-1)(a)]

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I am typing out the problem now it will be a few minutes sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[If h(x)=x+\sqrt{x} find h'(6)\]

OpenStudy (anonymous):

Wow that turned out bad can you read it?

OpenStudy (anonymous):

find the derivative of h(x)...plug in value of 6 for x what did you get for the derivative

OpenStudy (anonymous):

just use the rule for summations...

OpenStudy (anonymous):

1+1/2x^-1/2

OpenStudy (anonymous):

correct, now just switch x for 6

OpenStudy (anonymous):

Since its raised to the -1/2 it turns out really weird..

OpenStudy (anonymous):

that just means the square root is in the denominator of the fraction..

OpenStudy (anonymous):